OpenStudy (anonymous):

How many moles of boron are in 418.0 g of boron?

OpenStudy (julian.pitalua):

moles= mass/ atomic mass moles= 418.0g/10,81g/mol= xxxx mol

OpenStudy (nincompoop):

a proper dimensional analysis would be: \[grams_{substance} \times \frac{1 mole_{substance}}{grams_{substance}} = mole_{substance}\]

OpenStudy (nincompoop):

notice how grams cancel out?

OpenStudy (nincompoop):

if you're confused, that is because you do not know the proper definition of moles it is something you need to know first hand...